THE DATA DEPICTED BELOW IS FROM S. LTD. WHERE 300 EMPLOYEES WITH VARIOUS CATEGORIES OF PROFESSION ARE ON ROLLS.


Position Male female

Managerial 16 6

Professional 50 30

Technical 100 36

Clerical 18 44


Find the probability that:

1) A randomly selected employee is a female or a professional worker

2) A randomly selected employee is a female

3) A randomly selected employee who is found to hold a managerial position is a female

4) A randomly selected female employee having clerical position

5) A randomly selected employee given that he is a male is from technical position.


1
Expert's answer
2021-02-24T12:42:41-0500

We have that:




1) Probability that a randomly selected employee is a female or a professional worker:

P(female OR professional)=P(female)+P(professional)–P(female AND professional)=

"=\frac{116}{300}+\frac{80}{300}-\frac{30}{300}=\frac{166}{300}=0.55"

2) Probability that a randomly selected employee is a female:

"P(female)=\frac{116}{300}=0.39"

3) Probability that a randomly selected employee who is found to hold a managerial position is a female:

"P(female|managerial)=\frac{6}{22}=0.27"

4) Probability that a randomly selected female employee having clerical position:

"P(clerical|female)=\frac{44}{116}=0.38"

5) Probability that a randomly selected employee given that he is a male is from technical position:

"P(technical|male)=\frac{100}{184}=0.54"


Answer:

1) 0.55

2) 0.39

3) 0.27

4) 0.38

5) 0.54